### v2

# SSTP: Efficient Sample Selection for Trajectory Prediction

Abstract
Trajectory prediction is a core task in autonomous driving. However, training advanced trajectory prediction models on large-scale datasets is both time-consuming and computationally expensive. In addition, the imbalanced distribution of driving scenarios often biases models toward data-rich cases, limiting performance in safety-critical, data-scarce conditions. To address these challenges, we propose the Sample Selection for Trajectory Prediction (SSTP) framework, which constructs a compact yet balanced dataset for trajectory prediction. SSTP consists of two main stages (1) Extraction, in which a pretrained trajectory prediction model computes gradient vectors for each sample to capture their influence on parameter updates; and (2) Selection, where a submodular function is applied to greedily choose a representative subset that covers diverse driving scenarios. This approach significantly reduces the dataset size and mitigates scenario imbalance, without sacrificing prediction accuracy and even improving in high-density cases. We evaluate our proposed SSTP on the Argoverse 1 and Argoverse 2 benchmarks using a wide range of recent state-of-the-art models. Our experiments demonstrate that SSTP achieves comparable performance to full-dataset training using only half the data while delivering substantial improvements in high-density traffic scenes and significantly reducing training time. Importantly, SSTP exhibits strong generalization and robustness, and the selected subset is model-agnostic, offering a broadly applicable solution. The code is avaiable at https://github.com/RuiningYang/SSTP.
1 Introduction
Figure 1: Top: Comparison of model performance across varying scene densities. Superior enhancement on HiVT-64 [61] in high-density scenarios training with SSTP-selected 50% data from Argoverse 1 [5]. Bottom: Using the SSTP-selected subset results in a significant overall reduction in training time compared to using the full dataset, especially for complex models like HPNet [42].
Trajectory prediction aims to predict the future locations of agents conditioned on their past observations, which plays a key role in the domain of autonomous driving. This task is essential yet challenging due to the complex uncertain driving situations. With rapid developments in deep learning, various methods [61, 14, 4, 16, 52, 57, 30] have been proposed with promising trajectory prediction performance. Meanwhile, more large-scale realistic datasets [1, 49, 5, 47, 40] have been released by research institute and self-driving companies, which further push the boundary of this task. However, one common issue is that training these data-driven methods requires enormous computational resources and is time-consuming due to the large scale of the datasets. For example, the recent state-of-the-art method MTR [30] has over 66 million model parameters, and the Waymo Open Motion Dataset (WOMD) [40] has over 2.2 million trajectory samples. Training the complete model on this dataset requires a substantial amount of GPU hours, posing a significant computational burden. A similar situation arises in multiple methods as well as across various datasets, which is shown in Table 1. This phenomenon raises two following questions:
Q1: Do we need large-scale datasets to train the model?
Q2: How can we effectively reduce the training data volume without significantly compromising model accuracy?
Prompted by these questions, we take a deep dive into recent trajectory prediction benchmarks and reveal one finding the imbalance in the dataset. From a conventional machine learning perspective, standard tasks typically evaluate model performance using the average accuracy across all samples. However, in autonomous driving trajectory prediction, this metric inherently biases models toward driving scenarios with abundant data, leading to suboptimal performance in data-scarce scenarios. For safe driving, an ideal trajectory predictor should exhibit robust performance across both data-rich and data-scarce scenarios. However, due to the imbalanced distribution of driving scenarios, existing models struggle to maintain consistent performance across diverse scenarios, as illustrated in Figure 1.
Model Dataset Samples Training Time
HiVT [61] Argoverse 1 [5] 330k 43 hours
HPNet [42] Interaction [49] 62k 51 hours
HPNet [42] Argoverse 1 [5] 330k 256 hours
QCNet [62] Argoverse 2 [47] 250k 144 hours
DeMo [50] Argoverse 2 [47] 250k 72 hours
Table 1: Statistics on the training time of trajectory prediction models and the size of the datasets used. All training time data are derived from the original experimental setup and hardware specified by the model authors.
To address this issue, we introduce the Sample Selection for Trajectory Prediction (SSTP) framework, the first framework designed to construct a compact and balanced dataset for trajectory prediction. SSTP consists of two main stages: extraction and selection. In the extraction stage, a baseline model is pretrained on the full dataset for a few epochs to establish an initial feature representation. At the same time, the dataset is partitioned according to the number of agents per sample, and gradient vectors are computed for each sample to capture its influence on parameter updates. In the selection stage, we adopt a greedy strategy with a submodular function to compute scores and iteratively select representative samples. This approach builds a target subset that covers diverse driving scenarios while preserving the most distinctive features of each group. As shown in Figure 1, SSTP uniquely reduces data volume while maintaining, or even improving, model performance, particularly in high-density scenarios. Furthermore, it effectively balances the data distribution across various driving scenarios, making it well-suited for training a wide range of trajectory prediction models.
We evaluate our proposed method on the Argoverse 1 [5] and Argoverse 2 [47] datasets with multiple baseline methods. Empirical results demonstrate that SSTP successfully constructs a compact and well-balanced dataset. In high-density scenarios, training on the selected subset significantly outperforms using the full dataset, demonstrating the effectiveness of our method. Moreover, our work offers a resource-efficient dataset that maintains balanced performance across various driving scenarios, making it well-suited for training state-of-the-art trajectory prediction models. The main contributions are summarized as follows:
  • We introduce SSTP, a novel and effective framework for constructing a compact, balanced dataset for trajectory prediction, which boosts model performance across diverse driving scenarios.
  • SSTP significantly reduces training time and computational cost with a one-time selection process, making it efficient and adaptable to various baseline models.
  • Extensive experiments show that with only 50% of the original data, SSTP achieves comparable, or even better, performance across scenarios compared to full-dataset.
2 Related Work
2.1 Trajectory Prediction
Trajectory prediction infers an agent’s future motion from its historical observations. In recent years, research has increasingly concentrated on capturing complex multi-agent interactions, driving advances in various predictive methods [43, 15, 48, 25, 30, 61, 62, 42, 50, 23]. Furthermore, novel approaches including pretraining [7, 9, 21], historical prediction structure design [32, 42], GPT-style next-token prediction [34, 39], and post-processing optimization [60, 10] have significantly enhanced model performance, demonstrating strong results across various datasets. However, most state-of-the-art methods rely on training with large-scale datasets [1, 5, 40, 47, 13], leading to significant computational costs. In contrast, we propose a pioneering data selection strategy in the trajectory prediction domain that constructs a compact, balanced, yet highly representative dataset, significantly reducing training time while preserving model performance.
2.2 Long-Tail in Trajectory Prediction Dataset
The performance of the trajectory prediction models is evaluated based on the overall average. While they excel on benchmarks, these models often struggle with challengingscenarios [27, 46, 35] due to the long-tail data distribution, where common cases dominate and complex or rare situations are underrepresented [6]. Recent studies [46, 51, 58, 22] have started addressing the long-tail problem in trajectory prediction, primarily by leveraging contrastive learning to enhance feature representations. However, these methods mainly focus on optimizing feature-level learning while overlooking the distribution of data at the scenario level and the importance of individual samples. In contrast, our proposed method assesses the contribution of each sample and applies a refined selection strategy to build a balanced and compact dataset. Experimental results show that it significantly boosts performance in complex scenarios.
2.3 Training Sample Selection
Deep neural networks, especially Transformer-based models, depend on large-scale datasets and incur high computational costs. To reduce these costs and shorten training time, various methods focused on improving data efficiency have been proposed, including frequent parameter updates [37], fewer iterations [41], and dynamic learning rate adjustments [20, 12]. To directly reduce data volume, dataset condensation compresses raw data into compact synthetic samples [45, 55, 53, 19, 44, 54, 3]. Another widely studied approach is coreset selection, which constructs a weighted subset that closely approximates the statistical distribution of the original dataset [17, 11, 28, 29]. However, those approaches on dataset condensation [31, 26, 59, 38, 56] and coreset selection [18, 33] has been predominantly focused on image classification. In contrast, we propose a sample selection strategy based on submodular functions for the trajectory prediction domain. As a pioneering method, our approach significantly reduces the training data required while maintaining model performance.
3 Method
3.1 Problem Formulation
Consider a training set $\mathcal{D}=\{S_{j}\}_{j=1}^{M}$ consists of total $M$ driving samples. Each sample can be described by a triple $S=(X,Y,\mathcal{O})$, where $X$ denotes the observed trajectories, $Y$ represents the corresponding future trajectories for all the agents present, and $\mathcal{O}$ contains driving context information (e.g., maps). Trajectory prediction aims to estimate the future trajectory $Y$ conditioned on $X$ and $\mathcal{O}$. In this paper, our objective is to identify a smaller subset $\mathcal{C}\subseteq\mathcal{D}$ such that training a model in $\mathcal{C}$ stays within a limited budget $\mathcal{B}$. We define the budget as a fraction $\alpha$ of the full dataset size, i.e., $\mathcal{B}=\alpha|\mathcal{D}|$. At the same time, the model trained on $\mathcal{C}$ can achieve comparable trajectory prediction performance to that of the model trained on the full dataset $\mathcal{D}$.
Data Partitioning. As pointed out in Section 1, the data imbalance situation often occurs in commonly used datasets, specifically in large-scale ones. We can observe from Figure 2 that low-density scenes with few agents constitute a large portion of the dataset, while high-density scenes are underrepresented. This imbalance biases model training, making it more inclined to learn patterns from simpler scenarios while struggling to generalize to high-density scenarios. According to this observation, we first compute a density level $\rho(S_{j})$ for each sample $S_{j}\in\mathcal{D}$ based on the number of agents present. Then, we use a fixed interval $\tau$ to partition the dataset into $K$ disjoint subsets based on these density values. The partitioning can be defined as follows:
[…]
where $\rho_{min}$ is the minimum density level in $\mathcal{D}$.
Algorithm 1 Sample Selection for Trajectory Prediction
Input: Full Dataset $\mathcal{D}$, interval $\tau$, ratio $\alpha$, submodular function $P(\cdot)$
Output: Target dataset $\mathcal{C}$
1:Initialize: $\mathcal{C}\leftarrow\emptyset$;
2:Initialize: $\mathcal{B}\leftarrow\alpha|\mathcal{D}|$; $\triangleright$ Set budget
3:Partitioning: $\mathcal{D}_{k}$;
4:For $k\in\{K,K-1,...,1\}$: $\triangleright$ Reverse order
5:     $\mathcal{C}_{k}\leftarrow\emptyset$;
6:     $n_{k}\leftarrow\texttt{DynamicBudget}(\mathcal{B},k)$;
7:     if $n_{k}=|\mathcal{D}_{k}|$ then: $\triangleright$ Include all samples
8:         $\mathcal{C}\leftarrow\mathcal{C}\cup\mathcal{D}_{k}$;
9:     else:
10:         For $n=1$ to $n_{k}$: $\triangleright$ Iterate $n_{k}$ times
11:              $S_{j}\leftarrow\arg\min_{S_{j}\in\mathcal{D}_{k}\setminus\mathcal{C}_{k}}P(S_{j})$;
12:              $\mathcal{C}_{k}\leftarrow\mathcal{C}_{k}\cup\{S_{j}\}$;
13:         end for
14:         $\mathcal{C}\leftarrow\mathcal{C}\cup\mathcal{C}_{k}$;
15:     $\mathcal{B}\leftarrow\mathcal{B}-|\mathcal{C}_{k}|$; $\triangleright$ Update remaining budget
16:end for
17:Return $\mathcal{C}$
18:function DynamicBudget($\mathcal{B},k$):
19:     if $|\mathcal{D}_{k}|\leq\lfloor\mathcal{B}/{k}\rfloor$ then:
20:         $n_{k}\leftarrow|\mathcal{D}_{k}|$;
21:     else:
22:         $n_{k}\leftarrow\lfloor\mathcal{B}/{k}\rfloor$;
23:     Return $n_{k}$
Gradient Extraction. We calculate the total loss as follows:
$\mathcal{L}=\mathcal{L}_{reg}+\mathcal{L}_{cls},$ (2)
where $\mathcal{L}_{reg}$ measures the L2 norm difference between the best mode prediction and the ground truth trajectory, and $\mathcal{L}_{cls}$ aligns the mode predicted probabilities $\pi$ with the best mode.
By backpropagating the loss $\mathcal{L}$, we can obtain the gradient with respect to the model output $\hat{\mathbf{Y}}$ as follows:
$\nabla_{\hat{\mathbf{Y}}}\,\mathcal{L}=\nabla_{\hat{\mathbf{Y}}}\,(\mathcal{L}_{reg}+\mathcal{L}_{cls}).$ (3)
Finally, we perform an element-wise multiplication of the calculated gradient with the corresponding decoder latent vector $\mathbf{E}$ as follows:
$\mathbf{g}=\nabla_{\hat{\mathbf{Y}}}\,\mathcal{L}\odot\mathbf{E},$ (4)
where $\mathbf{g}$ captures the joint variations in the gradient and embedding spaces. In this manner, for each subset with $|\mathcal{D}_{k}|>n_{k}$, we construct a set of gradient feature vectors $\mathcal{G}_{k}=\{\mathbf{g}_{j}\}_{S_{j}\in\mathcal{D}_{k}}$ for all trajectory samples in that group.
Sample Selection. Given total $K$ disjoint subsets and the corresponding gradient feature vectors $\mathcal{G}_{k}$, we present a submodular gain function $P(\cdot)$ to evaluate the contribution of each sample. And we utilize a greedy algorithm to iteratively construct the target dataset $\mathcal{C}$ via optimizing the submodular gain function to ensure the selected target dataset optimally represents the entire dataset.
For each subset $\mathcal{D}_{k}$ with $|\mathcal{D}_{k}|>n_{k}$, where $n_{k}$ is the dynamic budget determined by the DynamicBudget function (see Algorithm 1), we initialize an empty set $\mathcal{C}_{k}$ to store the selected samples. For these subsets, let $\mathcal{G}_{k}$ denote the gradient feature vectors computed for the samples in $\mathcal{D}_{k}$. To evaluate the contribution of each sample, we define a submodular gain function based on a cosine similarity kernel. Specifically, for any trajectory sample $S_{j}\in\mathcal{D}_{k}$, the gain function is defined as follows:
$P(S_{j})=\sum_{S_{i}\in\mathcal{C}_{k}}\frac{\mathbf{g}_{i}\cdot\mathbf{g}_{j}}{\|\mathbf{g}_{i}\|\|\mathbf{g}_{j}\|}-\sum_{S_{i}\in\mathcal{D}_{k}\setminus\mathcal{C}_{k}}\frac{\mathbf{g}_{i}\cdot\mathbf{g}_{j}}{\|\mathbf{g}_{i}\|\|\mathbf{g}_{j}\|},$ (5)
where we use a cosine similarity kernel to measure the similarity between the sample $S_{j}$ and other samples. We then apply a greedy optimization strategy, iteratively selecting the sample follows:
$S^{*}=\arg\min_{S_{j}\in\mathcal{D}_{k}\setminus\mathcal{C}_{k}}P(S_{j}).$ (6)
At each iteration, the selected sample $S^{*}$ is added to $\mathcal{C}_{k}$. This process continues until the number of selected samples reaches the budget $n_{k}$ for subset $\mathcal{D}_{k}$. Note that we process the subsets starting with those having higher density levels, as these subsets tend to contain relatively fewer samples and are underrepresented. For subsets $\mathcal{D}_{k}$ where $|\mathcal{D}_{k}|\leq n_{k}$, we directly set $\mathcal{C}_{k}=\mathcal{D}_{k}$.
Finally, we can yield the target dataset $\mathcal{C}=\bigcup_{k=1}^{K}\mathcal{C}_{k}$. By incorporating gradient-based similarity into the submodular selection algorithm, our method ensures that the chosen subset maximally covers the gradient space while maintaining diversity, resulting in a smaller dataset that is both representative and informative, which ultimately leads to improved generalization in trajectory prediction models.
4 Experiments
Methods Ratio(%) HiVT-64 HiVT-128 HPNet
minADE minFDE xxMRxx minADE minFDE xxMRxx minADE minFDE xxMRxx
Argoverse 1 100 0.695 1.037 0.109 0.666 0.978 0.091 0.647 0.871 0.070
Random 60 0.745 1.163 0.132 0.719 1.078 0.129 0.680 0.951 0.091
Cluster 0.716 1.097 0.121 0.697 1.025 0.108 0.673 0.930 0.081
Herding 0.723 1.101 0.125 0.685 1.018 0.106 0.666 0.922 0.085
SSTP (Ours) 0.702 1.064 0.110 0.674 0.994 0.093 0.653 0.901 0.071
Random 50 0.750 1.175 0.137 0.728 1.098 0.126 0.687 0.967 0.091
Cluster 0.725 1.117 0.124 0.692 1.033 0.118 0.676 0.952 0.085
Herding 0.728 1.107 0.126 0.698 1.036 0.119 0.674 0.938 0.089
SSTP (Ours) 0.704 1.073 0.111 0.684 1.022 0.101 0.661 0.913 0.074
Random 40 0.752 1.183 0.139 0.727 1.109 0.126 0.696 0.987 0.099
Cluster 0.732 1.141 0.127 0.703 1.058 0.121 0.681 0.962 0.089
Herding 0.722 1.123 0.128 0.704 1.056 0.119 0.684 0.956 0.093
SSTP (Ours) 0.711 1.088 0.114 0.696 1.048 0.106 0.671 0.931 0.076
Table 2: Performance comparison results on Argoverse 1 [5] with data retention ratios of 60%, 50%, and 40%. The compared methods include Random Selection [36], K-Means Clustering [24], and Herding Selection [2, 8]. The model used for data selection is HiVT-64 [61], while the evaluation is conducted on HiVT-64, HiVT-128, and HPNet [42]. Ratio (%) represents the proportion of retained data relative to the full training set.
Method Ratio(%) QCNet DeMo
minADE minFDE xxMRxx minADE minFDE xxMRxx
Argoverse 2 100 0.724 1.258 0.162 0.657 1.254 0.163
Random 60 0.787 1.419 0.208 0.755 1.433 0.198
Cluster 0.773 1.406 0.192 0.693 1.386 0.187
SSTP (Ours) 0.740 1.316 0.163 0.682 1.344 0.164
Random 50 0.805 1.447 0.219 0.756 1.448 0.203
Cluster 50 0.798 1.435 0.193 0.732 1.437 0.191
SSTP (Ours) 0.754 1.352 0.172 0.704 1.414 0.173
Random 0.811 1.471 0.226 0.763 1.475 0.202
Cluster 40 0.813 1.495 0.214 0.732 1.456 0.195
SSTP (Ours) 0.778 1.410 0.183 0.723 1.450 0.191
Table 3: Performance comparison results on Argoverse 2 [47] with data retention ratios of 60%, 50%, and 40%. The compared methods include Random Selection and K-means clustering. The model used for data selection is QCNet [62], while the evaluation is conducted on QCNet and DeMo [50].
4.1 Benchmarks and Setup
Datasets. We evaluated the effectiveness of our proposed SSTP method on Argoverse Motion Forecasting Dataset 1.1  [5] and Argoverse 2 [47]. The Argoverse 1 dataset contains 323,557 real-world driving scenarios. All the training and validation scenarios are 5-second sequences sampled at 10 Hz. The length of the historical trajectory for each scenario is 2 seconds, and the length of the predicted future trajectory is 3 seconds. The Argoverse 2 dataset contains 250,000 scenarios, with the same sampling frequency of 10 Hz. Each trajectory has a larger observation window with 5 seconds and a longer prediction horizon with 6 seconds.
Baselines. For Argoverse 1, we validate our SSTP method on two SOTA models HiVT [61] and HPNet [42] for evaluation. For Argoverse 2, we evaluate our SSTP method using two SOTA models QCNet [62] and DeMo [50]. For a more comprehensive comparison, we also include three following data selection approaches:
(1) Random Selection [36]: randomly selects a certain proportion of training samples from the original training set.
(2) K-Means Clustering [24]: clusters trajectories within the observation window based on their features, and then selects the trajectory sample closest to the cluster center in each cluster as a representative.
(3) Herding Selection [2]: a greedy strategy that first computes the mean feature of all trajectories within the observation window and then iteratively selects trajectory samples that bring the mean of the selected subset as close as possible to the overall mean.
Metrics. Following the baselines, we also generate a total $6$ future trajectories and use the metrics minimum Average Displacement Error (minADE), minimum Final Displacement Error (minFDE), and Missing Rate (MR) to evaluate the prediction performance.
Implementation Details. We primarily utilized the pre-trained HiVT-64 and QCNet as backbone models to perform sample selection on the Argoverse 1 and Argoverse 2 datasets, respectively. To evaluate the performance of the selected subset, we follow their official training and validation protocols. We experimented with different selection ratios, different intervals, and assessed the prediction accuracy of the trajectory models after training on the corresponding subsets.
4.2 Main Results
Table 2 showcases the strong performance of our selected subset on the Argoverse 1 dataset across all compression rates. Following the same experimental setup as our baseline models, we trained HiVT and HPNet from scratch on the subset. When trained on the full dataset, HiVT-64 achieves a minADE of 0.695, minFDE of 1.037, and MR of 0.109. With a data retention rate of 60%, our selected subset achieves a minADE of 0.702, minFDE of 1.064, and MR of 0.110. This demonstrates that our method significantly reduces data volume while maintaining nearly lossless model performance. Even at a data retention rate of only 50%, models trained on our selected subset still maintain excellent performance, with minADE at 0.704, minFDE at 1.073, and MR at 0.111. Compared to random selection, our method reduces minADE by 0.046, minFDE by 0.102, and MR by 0.026. Moreover, while k-means clustering and herding methods show slight improvements in model performance, they still fall short of the performance achieved by our method. Furthermore, our subset also demonstrates superior performance on the HiVT128 model and the current SOTA model HPNet.
We further evaluated our proposed method on the Argoverse 2 dataset, which presents greater challenges due to its more diverse driving scenarios and longer prediction horizons. As shown in Table 3, our method consistently outperforms other data selection strategies across all data retention rates, achieving lower minADE and minFDE while maintaining a lower MR. These results further validate the robustness of our approach, as it maintains strong performance across different datasets. This demonstrates that our method is not only effective within a specific dataset but also generalizes well to more complex and diverse trajectory scenarios, such as those found in Argoverse 2.
Agent$<$40 Agent>=>=> =40 Agent>=>=> =60 Agent>=>=> =80
minADE minFDE xMRx minADE minFDE xMRx minADE minFDE xMRx minADE minFDE xMRx
Full 0.700 1.071 0.108 0.950 1.456 0.171 1.248 1.898 0.283 1.450 2.059 0.361
Random 0.734 1.127 0.119 0.997 1.552 0.193 1.287 1.989 0.315 1.638 2.359 0.389
SSTP (50%) 0.716 1.102 0.111 0.962 1.497 0.183 1.219 1.835 0.280 1.373 1.762 0.277
Table 4: Comparison of model performance across different scene densities when trained on the full dataset, random selection for 50% subset versus the 50% subset selected by our method (SSTP), where ours achieves superior performance.
Variants Selection Strategy Data Distribution(%) Model Performance
Partition Submodular Agent$<$40 Agent>=>=> =40 minADE minFDE xMRx
Full dataset - - 93.88 6.12 0.692 1.047 0.104
Random - - 85.16 14.84 0.741 1.164 0.125
SSTP w/ Submodular 93.88 6.12 0.724 1.115 0.116
SSTP w/ Partition 70.35 29.65 0.729 1.116 0.118
SSTP (Ours) 70.35 29.65 0.704 1.073 0.111
Table 5: Performance comparison of different data selection strategies on HiVT trained with Argoverse 1. This table illustrates the impact of Partition and selection with Submodular Gain strategies on data distribution and model performance. The whole dataset and random selection serve as baselines, while different variations of SSTP are evaluated. Our method (SSTP), which integrates both strategies, achieves the best results by maintaining a balanced data distribution and reducing minADE, minFDE, and MR.
4.3 Ablation Study
Performance Enhancement. Scene density in autonomous driving varies significantly. However, most existing trajectory prediction datasets predominantly focus on low-density scenarios. From the perspective of safe driving, an ideal trajectory predictor should maintain strong performance across diverse driving scenarios, regardless of data density. To address this, we conducted a comprehensive evaluation of our proposed method on existing models and across different scene densities. As presented in Table 4, our method consistently outperforms models trained on the full dataset, particularly in high-density agent scenarios, across both Argoverse 1 and Argoverse 2 datasets. When the agent density is below 40, our method achieves comparable performance to models trained on the full dataset, with only marginal increases of 0.01 in minADE and 0.03 in minFDE, while MR remains nearly unchanged. However, in high-density scenarios where the agent count exceeds 60, models trained on our selected subset demonstrate significant improvements, with minADE and minFDE errors decreasing by over 0.03 and 0.06, respectively. This advantage becomes even more pronounced in scenarios with more than 80 agents, where minADE is reduced by approximately 0.08, minFDE by nearly 0.3, and MR by nearly 9%. Similarly, the SOTA model HPNet, when trained on our selected subset, outperforms its counterpart trained on the full dataset, demonstrating consistent improvements on the Argoverse 2 dataset.
Density Balancing. Our method explicitly controls scene density distribution during selection, ensuring a more balanced dataset, as illustrated in Figure 2. In contrast, random selection fails to maintain this balance, resulting in an uneven distribution of scenarios with varying complexity, ultimately affecting the model’s generalization capability. As shown in Table 5 line 4, applying scene balancing alone already outperforms the random selection method. For instance, minADE decreases from 0.741 to 0.729, and minFDE drops from 1.164 to 1.116. This demonstrates that controlling scene density during data selection can enhance the effectiveness of trajectory prediction models. However, our full method outperforms scene balancing. By integrating submodular selection in addition to scene balancing, our approach further optimizes data selection by ensuring that the chosen samples carry higher information content. As a result, our method achieves minADE of 0.704 and minFDE of 1.073, further reducing errors. These findings indicate that while scene balancing is beneficial, it is insufficient to achieve optimal performance without also considering sample informativeness.
Effectiveness of Submodular Gain. To isolate the effect of submodular gain, we conducted an experiment where data selection was based solely on submodular importance scores, without considering scene density balancing, as shown in Table 5 line 3. The results indicate that using only submodular selection achieves a minADE of 0.724, lower than the 0.741 obtained through random selection, demonstrating that submodular-based sample selection improves data quality. However, it still underperforms compared to our full method. This is because prioritizing sample informativeness without adjusting for scene density leads to a dataset biased toward certain complexity levels, ultimately hindering the model’s generalization ability. In contrast, our full method, which integrates scene balancing with submodular gain, achieves the best performance across all metrics. These findings highlight the necessity of jointly considering both scene distribution balance and sample informativeness to construct an effective training dataset.
Figure 2: Top: Visualization examples of different driven scenarios in trajectory prediction dataset, using Argoverse 1 as an example. Bottom: Distribution of scene density before (left) and after (right) applying SSTP on Argoverse 1 with 50% retention ratio.
Data Retention Ratio. To examine the impact of different data retention ratios on model performance, we conducted experiments with retention rates of $\alpha$ = {60, 50, 40, 30, 20, 10}as shown in Figure 3. When higher model performance is required, retaining 50% of the data already achieves results comparable to training with on full dataset. Specifically, minADE and minFDE increase only slightly from 0.69 to 0.70, from 1.03 to 1.06, while the MR metric remains at a low level with 0.11. This demonstrates the effectiveness of our SSTP method, as the selected 50% subset is of higher quality compared to equally sized subsets selected by other methods. Furthermore, under limited computational resources, retaining only 20% of the data still yields reasonably good results. Although minADE and minFDE increase compared to the 50% subset (from 0.7063 to 0.7491 and from 1.0734 to 1.1666, respectively), the MR metric remains at 0.13. This indicates that even when the dataset size is reduced to $\frac{1}{5}$, our method can still maintain a reasonable level of predictive performance.
Epoch 0 5 8 10 15 64
minADE 0.713 0.704 0.707 0.708 0.710 0.712
minFDE 1.083 1.073 1.076 1.074 1.083 1.080
MR 0.112 0.111 0.111 0.111 0.111 0.111
Table 6: Performance comparison of models pretrained on the full dataset for different numbers of epochs. The pretrained weights are then used to initialize the model for SSTP, selecting 50% of the data.
Impact of Pretrained Backbone Epochs. To examine the influence of the pretrained backbone on subset selection, we conducted a series of experiments using models initialized identically but trained with different numbers of pretraining epochs. Taking HiVT-64 as an example, the official training setup involves training the model for 64 epochs using the full dataset. In our experiments, we varied the number of pretraining epochs as {0, 5, 8, 10, 15, 64} and analyzed its impact on subset selection, as shown in Table 6. The results indicate that moderate pretraining is crucial for effective subset selection. When the number of pretraining epochs is set to 5, the subset selection achieves optimal performance, consistently outperforming other configurations across all data retention ratios. As the pretraining epochs increase, subset selection continues to provide significant advantages over other data selection methods but does not surpass the performance observed at epoch 5. For models without pretraining, minADE degrades noticeably compared to models pretrained for 5 epochs. In contrast, at 64 pretraining epochs, as the model has already converged, the impact of sample selection on gradient updates diminishes. Although subset selection performance remains competitive, it does not yield further improvements over moderate pretraining.
4.4 Efficiency
Our method significantly reduces computational time while maintaining strong performance, as shown in Figure 1. On a single NVIDIA RTX 4090, training the HiVT model on the full dataset requires 7.78 hours, achieving a minADE of 0.69. In contrast, utilizing our SSTP method to select a 50% subset requires only 2.17 hours, significantly reducing the overall training time. When training on the selected subset, the total training time decreases to 6.37 hours (2.17 + 4.20), with only a minor increase of 0.01 in minADE. For the HPNet model, full dataset training takes 307.2 hours, whereas training with the selected subset reduces the training time by over 100 hours. This decisively confirms the superior efficiency of our approach in balancing training cost and model performance.
Figure 3: Performance of HiVT-64 and HPNet when training with subset chosen by different selection methods. Comparing at varying data retention ratios with minADE and minFDE. Methods include Random, K-means Clustering, Herding and SSTP, with the full dataset serving as a reference baseline. Lower values indicated better performance.
Backbone Ratio(%) HiVT-64 HPNet
minADE minFDE xMRx minADE minFDE xMRx
HiVT-64 50 0.704 1.073 0.111 0.661 0.913 0.074
60 0.703 1.065 0.111 0.654 0.901 0.072
HPNet 50 0.708 1.079 0.111 0.664 0.917 0.074
60 0.703 1.067 0.108 0.657 0.910 0.075
Table 7: Performance comparison of different backbone models, HiVT-64 and HPNet, trained on selected data at varying retention rates (50% and 60%). The table reports minADE, minFDE, and MR for both models under different data selection strategies.
4.5 Generalizability Study
Different Backbones. To further assess the generalizability of our method, we evaluated its performance using different backbone models. Specifically, we employed pre-trained HiVT-64 and HPNet as feature extractors on the Argoverse 1 dataset and conducted experiments at various data retention ratios. In Table 7, we report the results for the 50% and 60% retention ratios. Additional results for other retention ratios can be found in the Supplementary Material. As shown, regardless of the backbone used for subset selection, the final trajectory prediction performance remained nearly identical. These results suggest that our subset selection strategy is largely independent of the specific feature extractor, highlighting its robustness. This broad applicability makes it a promising approach for optimizing trajectory prediction models across diverse architectures.
Partition Interval. Given the number of agents in different trajectory prediction scenarios varies significantly, we examine the impact of scene density partition intervals on the selected subset to validate the generalizability of our method. Specifically, we divide the scenes in the Argoverse 1 dataset based on agent counts with partition intervals $\tau$ of {5, 10, 20}, forming multiple scene density categories. Within each category, we perform data selection. As shown in Table 8, the subsets selected using different partition intervals result in comparable model performance, with minimal variations in minADE and minFDE. This consistency across different settings highlights the robustness of our method. This suggests that our approach generalizes well to other datasets. When applying this method, the partition interval can be adjusted based on the dataset characteristics: if the dataset has relatively few agents per scene, a smaller interval is preferable; whereas for datasets with a high variance in agent count, a larger interval may be more suitable to better balance scene density.
Partition minADE minFDE MR
$\tau$ = 5 0.703 1.056 0.110
$\tau$ = 10 0.702 1.064 0.111
$\tau$ = 20 0.707 1.081 0.113
Table 8: Comparison of data selected by SSTP with different partition intervals $\tau$ and its impact on HiVT performance using the Argoverse 1 dataset.
5 Conclusion
In this paper, we presented the Sample Selection for Trajectory Prediction (SSTP) framework, a novel, data-centric approach that constructs a compact yet balanced dataset for trajectory prediction. SSTP effectively tackles the challenges posed by data imbalance and the high training costs inherent in large-scale trajectory datasets. By significantly reducing the training data volume while maintaining, and even enhancing the model performance in high-density scenarios. SSTP not only accelerates training but also delivers results comparable to, or better than, those achieved with full-dataset training. Extensive evaluations on the Argoverse 1 and Argoverse 2 benchmarks across a wide range of state-of-the-art models underscore the practical value of our approach in improving both efficiency and robustness in trajectory prediction for autonomous driving.
Limitations. Despite the effectiveness of our proposed method, improvements could be made to further reduce the computation overhead of our sample selection process. Further optimization can be made to streamline this process. Additionally, investigating methods to sustain model performance under extremely low data retention rates, e.g., 10%, presents a promising path for future research.